OpenStudy (anonymous):

The coordinates J(2, 2), K(-1, 3), L(-2, -1) form a right triangle. true or false

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Plot the points. Then use the distance formula. If the Pythagorean theorem holds, then it's a right triangle.

OpenStudy (anonymous):

i dont have anything to plot them on

OpenStudy (anonymous):

\[d=\sqrt{\Delta x ^{2}+\Delta y ^{2}}\]

jimthompson5910 (jim_thompson5910):

you need to use the distance formula to find the segments JK KL JL

jimthompson5910 (jim_thompson5910):

if you can show that a^2 + b^2 = c^2 is true, where a,b,c are the 3 side lengths and c is the longest side, then the triangle is a right triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

so its kinda long process to find im guessing okayyyyyyy here we go

jimthompson5910 (jim_thompson5910):

yes it's quite long since you need to use the formula hairynmad gave you to find JK, KL and JL

jimthompson5910 (jim_thompson5910):

once you have those 3 sides, you plug them into a^2 + b^2 = c^2 to test if it's a right triangle or not

OpenStudy (anonymous):

can you help e set up the first one just to make sure im doing it right

jimthompson5910 (jim_thompson5910):

which side length do you want to find JK KL JL ??

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

is the answer to it sqrt10

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so after i get all those answers how do i plug them into the a2+b2=c2?

OpenStudy (anonymous):

like does it matter where i place them in the equation?

OpenStudy (dwelsher):

false

OpenStudy (anonymous):

longest side will be c, the hypotenuse the other two, the legs, don't matter, can be either a or b

jimthompson5910 (jim_thompson5910):

as long as c is the longest side, everything else doesn't matter

OpenStudy (anonymous):

okay gotcha . so is it sqrt10 or no?

jimthompson5910 (jim_thompson5910):

correct, JK = sqrt(10)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

sooooo this is what i got

OpenStudy (anonymous):

sqrt10^2+ sqrt17 ^2 = sqrt25^2

OpenStudy (anonymous):

hwo the heck do i work that out

jimthompson5910 (jim_thompson5910):

one sec while I check

OpenStudy (anonymous):

i got false it is not a right triangle am i right?

OpenStudy (anonymous):

okay

jimthompson5910 (jim_thompson5910):

KL = sqrt(17) that's correct

jimthompson5910 (jim_thompson5910):

JL = sqrt(25) which is really just 5 so JL = 5

jimthompson5910 (jim_thompson5910):

a^2 + b^2 = c^2 (JK)^2 + (KL)^2 = (JL)^2 (sqrt(10))^2 + (sqrt(17))^2 = (5)^2 10 + 17 = 25 27 = 25 This is FALSE, so a^2 + b^2 = c^2 is FALSE when a = sqrt(10) b = sqrt(17) c = 5

jimthompson5910 (jim_thompson5910):

so yes, you're correct this isn't a right triangle

OpenStudy (anonymous):

okay wow that really helps i appreciate the help both of you so much

jimthompson5910 (jim_thompson5910):

you're welcome